# Text-Trained LLMs Can Zero-Shot Extrapolate PDE Dynamics

### v1

@@ A.5 Results for Other Model Families
We replicate the full experimental setup from Section 4 using models outside the Llama family: Phi-4-14B and SmolLM-3-3B, with Llama-3.2-3B (the representative model analyzed in the main text) included for comparison. Figures 15 and 16 summarize the results.
@@
As shown in Figure 15, the overall qualitative trends remain consistent with those reported in the main text. One-step prediction accuracy improves with longer input context and degrades with increasing output length, while multi-step rollouts exhibit algebraic error accumulation. For models with similar parameter counts—e.g., Llama-3.2-3B and SmolLM-3-3B—the exact quantitative prediction errors differ slightly. This reinforces the observation from the main text that the “model size effect” we report arises primarily when comparing models within the same family; models from different families with similar parameter sizes can exhibit slightly different prediction errors, likely reflecting differences in architecture, training data, and other design choices.
@@
Similarly, Figure 16 shows that the prediction uncertainty trends are consistent with those observed for Llama-3 models. Specifically, spatial entropy progresses through three distinct learning stages as input context increases, grows with output length, and separator tokens (e.g., commas) are predicted with near-perfect confidence across all settings. The main deviation is that Phi‑4‑14B consistently exhibits higher mean spatial entropy than the two 3B models. This difference is largely attributable to the inference temperature: Llama-3777https://huggingface.co/meta-llama/Llama-3.2-3B/tree/main and SmolLM-3888https://huggingface.co/HuggingFaceTB/SmolLM3-3B/tree/main use $T=0.6$, the default generation temperature specified in their configuration files, while Phi-4999https://huggingface.co/microsoft/phi-4/tree/main defaults to $T=1.0$, as its configuration omits a temperature setting. Lower temperatures ($T<1$) scale the logits to increase their relative magnitude, yielding sharper (lower-entropy) output distributions. In contrast, higher temperatures ($T>1$) reduce these relative differences, producing flatter (higher-entropy) distributions. The temperature-scaled softmax function (Goodfellow et al., 2016) is defined as:
@@
Figure 15: Prediction accuracy of Phi-4-14B and SmolLM-3-3B models (with Llama-3.2-3B for reference), using the same experimental setup described in Sections 4.1 and 4.2.
Figure 16: Uncertainty analysis of Phi-4-14B and SmolLM-3-3B models, with Llama-3.2-3B included for reference, using the same experimental setup described in Section 4.3 and Figure 5. Top: Mean spatial entropy $\bar{H}$ as a function of context length $N_{\mathrm{T}}$ (left) and output length $N_{\mathrm{X}}$ (right). Bottom: Token-level softmax distributions at the syntax-only stage for Phi-4-14B. The overall entropy behavior remains consistent with the Llama-3 models, with one notable difference: Phi-4-14B exhibits higher average spatial entropy than the two 3B models shown here, in nearly all cases, due to its higher default inference temperature (see Appendix A.5 for discussion). A pronounced peak in average spatial entropy also appears at very short contexts ($N_{\mathrm{T}}=2$), similar to Llama-3.1-8B-Instruct, where increased uncertainty arises from spatial value tokens more frequently acting as generic placeholders rather than producing deterministically incorrect outputs.
A.6 Architectural Details and Size Comparison of Llama-3 Models
Table 1 summarizes the architectural configurations of the three Llama-3 models evaluated in our main experiments. All values are sourced directly from Meta’s official generation_config files, released on Hugging Face. [^10] https://huggingface.co/meta-llama Differences in parameters such as hidden size, number of layers, and attention configurations account for the varying model sizes, which in turn influence the emergence of ICL behaviors observed in zero-shot solutions of PDEs.
Table 1: Architectural details of evaluated Llama-3 models
@@ A.6 Architectural Details and Size Comparison of Llama-3 Models
A.7 Validation of Non-Trivial Temporal Evolution
@@
To empirically validate that our prediction tasks (Section 3) are non-trivial, we analyze the temporal differences $Q_{i,j+1}-Q_{i,j}$ across the spatial grid under the finest discretization setting ($N_{\mathrm{X}}=40$, $N_{\mathrm{T}}=50$). As shown in Figure 17, the discretized solution exhibits meaningful variation between adjacent time steps across the spatial domain, confirming that the system evolves in a non-trivial manner over time. Coarser discretizations (e.g., $N_{\mathrm{X}}=14$, $N_{\mathrm{T}}=25$), as used in the multi-step rollout task, naturally introduce larger changes between time steps due to increased temporal spacing compared to the finest setting shown here. These observations support the discretization design choices used in the main experiments, which preserve clear spatiotemporal variation and ensure that model performance reflects an understanding of the underlying PDE dynamics rather than relying on trivial extrapolation strategies.
@@
Figure 17: Temporal differences $Q_{i,j+1}-Q_{i,j}$ at each spatial grid point for the finest discretization setting ($N_{\mathrm{X}}=40,N_{\mathrm{T}}=50$) used in the experimental setups of Section 4, where ${\bm{Q}}\in\mathcal{Z}^{N_{\mathrm{X}}\times(N_{\mathrm{T}}+1)}$ is the quantized representation of the PDE solution used as input to the LLM (see Section 3). The heatmap shows that, even at this resolution, the discretized solution exhibits meaningful changes between adjacent time steps, indicating that the prediction task requires modeling nontrivial temporal evolution. Since coarser discretizations correspond to larger time steps, they naturally induce greater local variation, further supporting the non-trivial nature of the prediction task across discretizations.
A.8 Additional Multi-Step Rollout Visualizations
@@
Figure 18: Multi-step prediction and error visualization for a different randomly sampled initial condition, using the same experimental setup as in Section 4.2.
@@
Figure 18 presents an additional multi-step rollout using a different initial condition, sampled with np.random.seed(42) following the procedure in Appendix A.1. The main text example (Figure 3) uses np.random.seed(1). Changing the seed alters the sampled interior values, yielding distinct spatiotemporal trajectories. The results here corroborate those in Section 4.2, indicating that the observed model behaviors generalize across initial conditions.
A.9 Error Patterns and Capacity Limitations in the Llama-3.2-1B Model
@@
Figure 19: Bias analysis of the multi-step prediction rollout from Figure 3 for the Llama-3.2-1B model. Left: Absolute error vs. reconstructed prediction value shows no correlation, indicating that prediction errors are not biased toward particular output magnitudes or special integer-like values (e.g., 200, 300, 400), and suggesting that tokenization procedures are unlikely to be the source of the error. Center & Right: Absolute error vs. spatial and temporal derivatives. In this rollout, higher errors cluster in regions with low local variation, though this trend does not persist across other random initial conditions.
@@
In this appendix, we analyze systematic errors exhibited by the smallest model, Llama-3.2-1B, during multi-step PDE rollouts over the random initial condition analyzed in Section 4.2. As shown in Figure 3, the 1B model produces structured errors that grow over the 10-step prediction horizon and consistently concentrate at specific spatial locations, in contrast to the low-magnitude errors from the 3B and 8B models, which remain more evenly distributed across the spatial domain. These persistent error patterns from the 1B model, averaged over 20 generations from the same initial condition, suggest a systematic prediction bias likely attributable to the model’s inductive biases or capacity limitations, rather than to stochasticity introduced during LLM sampling at inference time.
@@
Figure 19 investigates the origin of this systematic prediction bias by analyzing how the prediction errors relate to properties of the discretized PDE solution—specifically, its magnitude and local variation in space and time. The left panel shows errors grouped by the reconstructed floating-point values of the 1B model’s outputs. The lack of any discernible trend suggests that the bias is not driven by output magnitude or tokenization artifacts such as a preference for special integer-like values (e.g., 200, 300, 400). The center and right panels display errors grouped by spatial and temporal derivatives, respectively, both approximated using finite-difference stencils (central differences for interior points and forward/backward differences at boundaries). In this specific rollout, higher errors tend to occur in regions with lower local variation, hinting at a potential trend. However, this behavior is not consistent across different initial conditions.
@@
A.10 Accumulation of Predictive Uncertainty in Multi-Step Rollouts
Figure 20: Mean spatial entropy $\bar{H}$ as a function of prediction step $n$, using the multi-step rollout setup from Section 4.2 and the uncertainty metric defined in Section 4.3. Shaded regions denote 95% confidence intervals across 50 randomly sampled initial conditions, with one LLM rollout per condition.
In this appendix, we extend the entropy-based uncertainty analysis described in Section 4.3 for one-step predictions to the multi-step rollout setup introduced in Section 4.2, where the LLM autoregressively generates future time steps by appending its own outputs as additional inputs at each step. Since no natural language prompting is used, the model does not distinguish between ground-truth context and its own generated predictions, allowing us to directly analyze its ICL capacity to roll out PDE dynamics purely from serialized numerical input.
Under fixed spatial and temporal discretization, all models exhibit a consistent decrease in mean spatial entropy $\bar{H}$ with increasing prediction horizon (Figure 20), reflecting progressively more deterministic outputs. Notably, the smaller model (Llama-3.2-1B) maintains substantially higher entropy throughout the rollout compared to larger models (Llama-3.1-8B and Llama-3.2-3B), indicating that the model size influences the confidence level of multi-step predictions. However, this trend toward increased confidence does not correspond to improved predictive accuracy: as shown in Figure 4, errors grow algebraically over time. This illustrates that prediction uncertainty may decrease due to internal belief reinforcement within the LLM, even as predictive accuracy degrades.
A.11 Token-Level Distribution Visualizations Across Learning Stages
To complement the stage-wise analysis in Section 4.3, Figure 21 visualizes token-level softmax distributions from Llama-3.1-8B across all three ICL stages—syntax-only, exploratory, and consolidation—for a representative initial condition (same as in the main text). The syntax-only stage is shown in the bottom panel of Figure 5; this appendix provides a more comprehensive view by including the later stages as well. Together, these illustrate how the model’s predictive uncertainty evolves with increasing context length: initially focused on reproducing surface-level syntax, then entering a phase of high uncertainty as it explores plausible continuations, and ultimately converging to confident predictions that align with the underlying spatiotemporal PDE dynamics.

### v2

@@ 5 Results for Other Model Families
We replicate the full experimental setup from the Results section of the main text using models outside the Llama family: Phi-4-14B and SmolLM-3-3B, with Llama-3.2-3B (the representative model analyzed in the main text) included for comparison. Figures S10 and S11 summarize the results.
@@
As shown in Fig. S10, the overall qualitative trends remain consistent with those reported in the main text. One-step prediction accuracy improves with longer input context and degrades with increasing output length, while multi-step rollouts exhibit algebraic error accumulation. For models with similar parameter counts—e.g., Llama-3.2-3B and SmolLM-3-3B—the exact quantitative prediction errors differ slightly. This reinforces the observation from the main text that the “model size effect” we report arises primarily when comparing models within the same family; models from different families with similar parameter sizes can exhibit slightly different prediction errors, likely reflecting differences in architecture, training data, and other design choices.
@@
Similarly, Fig. S11 shows that the prediction uncertainty trends are consistent with those observed for Llama-3 models. Specifically, spatial entropy progresses through three distinct learning stages as input context increases, grows with output length, and separator tokens (e.g., commas) are predicted with near-perfect confidence across all settings. The main deviation is that Phi‑4‑14B consistently exhibits higher mean spatial entropy than the two 3B models. This difference is largely attributable to the inference temperature: Llama-3 [44] and SmolLM-3 [45] use $T=0.6$, the default generation temperature specified in their configuration files, while Phi-4 [46] defaults to $T=1.0$, as its configuration omits a temperature setting. Lower temperatures ($T<1$) scale the logits to increase their relative magnitude, yielding sharper (lower-entropy) output distributions. In contrast, higher temperatures ($T>1$) reduce these relative differences, producing flatter (higher-entropy) distributions. The temperature-scaled softmax function [47] is defined as:
@@
Figure S10: Prediction accuracy of Phi-4-14B and SmolLM-3-3B models (with Llama-3.2-3B for reference), using the same experimental setup described in the Results section of the main text.
Figure S11: Uncertainty analysis of Phi-4-14B and SmolLM-3-3B models, with Llama-3.2-3B included for reference, using the same experimental setup described in the Results section of the main text. Top: Mean spatial entropy $\bar{H}$ as a function of context length $N_{\mathrm{T}}$ (left) and output length $N_{\mathrm{X}}$ (right). Bottom: Token-level softmax distributions at the syntax-only stage for Phi-4-14B. The overall entropy behavior remains consistent with the Llama-3 models, with one notable difference: Phi-4-14B exhibits higher average spatial entropy than the two 3B models shown here, in nearly all cases, due to its higher default inference temperature (see Section 5 for discussion). A pronounced peak in average spatial entropy also appears at very short contexts ($N_{\mathrm{T}}=2$), similar to Llama-3.1-8B-Instruct, where increased uncertainty arises from spatial value tokens more frequently acting as generic placeholders rather than producing deterministically incorrect outputs.
6 Architectural Details and Size Comparison of Llama-3 Models
Table S1 summarizes the architectural configurations of the three Llama-3 models evaluated in our main experiments. All values are sourced directly from Meta’s official generation_config files, released on Hugging Face [48, 44, 49]. Differences in parameters such as hidden size, number of layers, and attention configurations account for the varying model sizes, which in turn influence the emergence of ICL behaviors observed in zero-shot solutions of PDEs.
Table S1: Architectural details of evaluated Llama-3 models
@@ 6 Architectural Details and Size Comparison of Llama-3 Models
7 Validation of Non-Trivial Temporal Evolution
@@
To empirically validate that our prediction tasks are non-trivial, we analyze the temporal differences $Q_{i,j+1}-Q_{i,j}$ across the spatial grid under the finest discretization setting ($N_{\mathrm{X}}=40$, $N_{\mathrm{T}}=50$). As shown in Fig. S12, the discretized solution exhibits meaningful variation between adjacent time steps across the spatial domain, confirming that the system evolves in a non-trivial manner over time. Coarser discretizations (e.g., $N_{\mathrm{X}}=14$, $N_{\mathrm{T}}=25$), as used in the multi-step rollout task, naturally introduce larger changes between time steps due to increased temporal spacing compared to the finest setting shown here.
In addition to confirming that the extrapolation task remains non-trivial even under the finest discretization, we further compare LLM one-step predictions against two repeat-based baselines: a temporal-repeat baseline that uses the final in-context time slice as the predicted time slice, and a last-token-repeat baseline that fills the predicted time slice with the final scalar token of the input. As shown in Fig. S13 (left), once sufficient temporal context is provided, all Llama-3 models outperform both baselines, with the Llama-3.1-8B and Llama-3.2-3B achieving errors roughly one order of magnitude below the temporal-repeat baseline and two to three orders of magnitude below the last-token-repeat baseline. This difference indicates that the models are learning the underlying spatiotemporal dynamics rather than relying on naive continuation strategies. The right panel of Fig. S13 fixes the temporal context at $N_{\mathrm{T}}=50$ and varies the spatial resolution $N_{\mathrm{X}}$. In this finely time-discretized regime, the temporal-repeat baseline itself becomes a non-trivial predictor. While the Llama-3.1-8B and Llama-3.2-3B continue to outperform both baselines across resolutions, the Llama-3.2-1B degrades rapidly and eventually falls below the temporal-repeat baseline. This behavior reflects the limited in-context learning capacity of the smallest model at finer spatial discretizations, consistent with the scaling trends discussed in the main text.
These observations support the discretization design choices used in the experiments, which preserve clear spatiotemporal variation and ensure that model performance reflects an understanding of the underlying PDE dynamics rather than relying on naive extrapolation strategies. A corresponding evaluation for multi-step rollouts, where the extrapolation challenge becomes more pronounced, is provided in Section 8.3. The analogous comparison reported there further strengthens these conclusions.
@@
Figure S12: Temporal differences $Q_{i,j+1}-Q_{i,j}$ at each spatial grid point for the finest discretization setting ($N_{\mathrm{X}}=40,N_{\mathrm{T}}=50$) used in the experimental setups of the Results section of the main text, where $\bm{Q}\in\mathcal{Z}^{N_{\mathrm{X}}\times(N_{\mathrm{T}}+1)}$ is the quantized representation of the PDE solution used as input to the LLM (see Methods section of the main text). The heatmap shows that, even at this resolution, the discretized solution exhibits meaningful changes between adjacent time steps, indicating that the prediction task requires modeling nontrivial temporal evolution. Since coarser discretizations correspond to larger time steps, they naturally induce greater local variation, further supporting the non-trivial nature of the prediction task across discretizations.
Figure S13: LLM performance compared to repeat-based baselines. The temporal-repeat baseline uses the final in-context time slice as the prediction for the next time slice, and the last-token baseline fills the next time slice with the final scalar token of the input. Left: With sufficient temporal context, all Llama-3 models outperform both baselines, indicating genuinely non-trivial temporal extrapolation. Right: Under fine temporal discretization ($N_{\mathrm{T}}=50$), the temporal-repeat baseline provides a non-trivial continuation of the solution. The 8B and 3B models continue to outperform both baselines across spatial resolutions, while the 1B model degrades with increasing $N_{\mathrm{X}}$ and eventually falls below the temporal-repeat baseline, reflecting its limited capacity at finer spatial discretizations.
8 Additional Multi-Step Rollout Results
@@
Figure S14: Multi-step prediction and error visualization for the Allen–Cahn equation with a different randomly sampled initial condition, using the same experimental setup as in the Results section of the main text.
Figure S15: Detailed visualizations corresponding to Fig. 3 of the main text, with additional numerical benchmarks and results from the Llama-3.2-3B model.
8.1 Additional Multi-Step Rollout Visualizations
@@
Figure S14 presents the multi-step prediction error trend for the Allen–Cahn equation with a different initial condition, sampled with np.random.seed(42) following the procedure in Section 1. Figure 3 of the main text uses np.random.seed(1) for Allen–Cahn and np.random.seed(42) for wave. Changing the seed alters the sampled interior values, yielding distinct spatiotemporal trajectories. The results here corroborate those in the Results section of the main text, indicating that the observed model behaviors generalize across initial conditions.
Figure S15 provides the detailed visualizations corresponding to Fig. 3 of the main text. It shows representative multi-step rollouts for the Allen–Cahn and wave equations with one additional numerical benchmark each: IMEX for Allen–Cahn and Crank–Nicolson for the wave equation. This figure also includes results for the Llama-3.2-3B model, which were omitted from the main text due to space constraints. The extended visualizations further illustrate that the smallest model (Llama-3.2-1B) struggles to sustain coherent PDE dynamics over extended horizons, while numerical benchmarks confirm consistency with classical solvers.
8.2 Multi-Step Rollouts on Non-Uniform Grids
We additionally evaluate the models on solutions sampled on a non-uniform Chebyshev spatial grid that clusters near the boundaries [50], while keeping all other experimental settings identical to the multi-step prediction setup described in the Results section of the main text. Specifically, the Chebyshev points, labeled in decreasing order from $x_{0}=1$ to $x_{N}=-1$, are given by $x_{k}=\cos(k\pi/N),k=0,1,\dots,N$. Figure S16 shows that the qualitative behaviors of all models remain consistent with the uniform-grid experiments: errors accumulate algebraically with the rollout horizon. These results indicate that the observed trends are not limited to uniform spatial grids.
8.3 Multi-Step Prediction Performance Against Naive Temporal Baselines
We compare the LLM rollouts against two naive temporal baselines: temporal repeat and a linear autoregressive model with one-step memory (AR1), in the multi-step prediction setup described in the Results section for the Allen–Cahn equation. We use a longer simulation time $T=1$ with the same spatial and temporal discretization as in the main text, giving each method access to $31$ context steps (including the initial condition), $\{u(x_{i},t_{j})\}_{i=1,j=0}^{N_{\mathrm{X}},30}$. The LLM then autoregressively generates $15$ prediction steps without access to intermediate ground truth. The temporal-repeat baseline copies the final in-context time slice for all future steps. The AR1 baseline uses a linear model trained on the same temporal context to map a state $u(\cdot,t_{j})$ to the subsequent state $u(\cdot,t_{j+1})$ by minimizing the residual sum of squares between predicted and observed next-time-slice values; it is then rolled out autoregressively. As shown in Fig. S17, the LLMs consistently achieve substantially lower multi-step prediction errors across longer rollout horizons, indicating that they capture and propagate the underlying PDE dynamics rather than relying on naive continuation methods.
Figure S16: Multi-step prediction errors on uniform and Chebyshev spatial grids under the same experimental setup as in the Results section of the main text, showing that the observed error-growth trends are not limited to uniform discretizations.
Figure S17: Multi-step prediction errors for LLM rollouts compared with two naive temporal baselines: temporal repeat and a one-step linear autoregressive model (AR1). All methods receive 31 in-context time slices and autoregressively predict 15 future steps for the Allen–Cahn equation over a simulation horizon of $T=1$, using the same discretization and experimental settings as in the Results section of the main text. The LLMs achieve substantially lower errors than both baselines across longer rollout horizons, demonstrating that their predictions reflect the underlying PDE dynamics rather than naive temporal continuation.
9 Error Patterns and Capacity Limitations in the Llama-3.2-1B Model
@@ 8.3 Multi-Step Prediction Performance Against Naive Temporal Baselines
Figure S18: Multi-step rollouts and error analysis for the Llama-3.2-1B model. Top: Predictions averaged over 20 LLM repeats for the randomly sampled initial conditions used in the main text. Bottom: Bias analysis. Left: Absolute error vs. reconstructed prediction value shows no correlation, indicating that errors are not biased toward particular magnitudes or integer-like values (e.g., 200, 300, 400), suggesting tokenization procedures are unlikely to be the source. Center and right: Errors vs. spatial and temporal derivatives. In this rollout, larger errors cluster in regions of low variation, though this trend does not persist across other initial conditions.
@@
In this section, we analyze systematic errors exhibited by the smallest model, Llama-3.2-1B, during multi-step PDE rollouts. Figure S18 shows predictions averaged over 20 LLM repeats for the randomly sampled initial condition in the main text alongside bias analyses of the resulting errors. The 1B model produces structured errors that grow over the 10-step prediction horizon and consistently concentrate at specific spatial locations, in contrast to the low-magnitude errors from the 3B and 8B models, which remain more evenly distributed across the spatial domain. These persistent error patterns from the 1B model, averaged over 20 repeats from the same initial condition, suggest a systematic prediction bias likely attributable to the model’s inductive biases or capacity limitations, rather than to stochasticity introduced during LLM sampling at inference time.
@@
The bottom panels of Fig. S18 investigate the origin of this systematic prediction bias by analyzing how the prediction errors relate to properties of the discretized PDE solution—specifically, its magnitude and local variation in space and time. The left panel shows errors grouped by the reconstructed floating-point values of the 1B model’s outputs. The lack of any discernible trend suggests that the bias is not driven by output magnitude or tokenization artifacts such as a preference for special integer-like values (e.g., 200, 300, 400). The center and right panels display errors grouped by spatial and temporal derivatives, respectively, both approximated using finite-difference stencils (central differences for interior points and forward/backward differences at boundaries). In this specific rollout, higher errors tend to occur in regions with lower local variation, hinting at a potential trend. However, this behavior is not consistent across different initial conditions.
@@
10 Accumulation of Predictive Uncertainty in Multi-Step Rollouts
Figure S19: Mean spatial entropy $\bar{H}$ as a function of prediction step $n$, using the multi-step rollout setup and the uncertainty metric defined in the Results section of the main text. Shaded regions denote 95% confidence intervals across 50 randomly sampled initial conditions, with one LLM rollout for each initial condition.
In this section, we extend the entropy-based uncertainty analysis for one-step predictions to the multi-step rollout setup described in the Results section of the main text, where the LLM autoregressively generates future time steps by appending its own outputs as additional inputs at each step. Since no natural language prompting is used, the model does not distinguish between ground-truth context and its own generated predictions, allowing us to directly analyze its ICL capacity to roll out PDE dynamics purely from serialized numerical input.
Under fixed spatial and temporal discretization, all models exhibit a consistent decrease in mean spatial entropy $\bar{H}$ with increasing prediction horizon (Fig. S19), reflecting progressively more deterministic outputs. Notably, the smaller model (Llama-3.2-1B) maintains substantially higher entropy throughout the rollout compared to larger models (Llama-3.1-8B and Llama-3.2-3B), indicating that the model size influences the confidence level of multi-step predictions. However, this trend toward increased confidence does not correspond to improved predictive accuracy: as shown in Fig. 4 of the main text, errors grow algebraically over time. This illustrates that prediction uncertainty may decrease due to internal belief reinforcement within the LLM, even as predictive accuracy degrades.
11 Token-Level Distribution Visualizations Across Learning Stages
To complement the stage-wise analysis in the Results section of the main text, Fig. S20 visualizes token-level softmax distributions from Llama-3.1-8B across all three ICL stages—syntax-only, exploratory, and consolidation—for a representative initial condition (same as in the main text). Figure 5(c) of the main text shows these distributions only at odd spatial positions; this section provides the complete set across all spatial positions. Together, these illustrate how the model’s predictive uncertainty evolves with increasing context length: initially focused on reproducing surface-level syntax, then entering a phase of high uncertainty as it explores plausible continuations, and ultimately converging to confident predictions that align with the underlying spatiotemporal PDE dynamics.